OpenStudy (anonymous):

Fe2O3+2Al=Al2O3+2Fe A welder has 1.873 × 10^2 g Fe2O3 and 94.51 g Al in his welding kit. Which reactant will he run out of first? (options: Fe2O3 or Al.......it's Fe2O3). How much of this reactant should he order to make sure he runs out of both reactants at the same time? He should order ___g of the limiting reactant.

OpenStudy (jhannybean):

irst convert the number of grams of each to moles. \[\sf 1.873 ~x~ 10^2 \text{g}~ Fe_2O_3\cdot \frac{1~\text{mol}~Fe_2O_3}{159.7~\text{g}~Fe_2O_3} =~?~ mol~Fe_2O_3 ~(have) \]\[\sf 94.51~g~Al \cdot \frac{1 ~mol ~Al}{26.98~g ~Al}=~? ~mol~Al~(have)\]Then pick one of the two to do a mole to mole ratio. \[\sf ~?~mol~Al \cdot \frac{1~mol~Fe_2O_3}{2~mol~Al}=~?~\text{mol of}~~Fe_2O_2 ~is~needed~\]If the amount needed is bigger than the amount you have, then the element or compound you used to find the mole ratio is the LR. If the moles of \(\sf Fe_2O_3\) needed is less than what you have, then that element or compound is the excess.

OpenStudy (anonymous):

Thank you soooo much!!!!